# references.bib

### v2

@@ Conclusions
We developed PHEONA, an evaluation framework for applying LLM-based methods to computational phenotyping tasks. This framework can be used to assess whether LLMs are feasible and if so, which model is best for the task of interest. We provided an example use case to demonstrate the framework and promote further research into improving traditional computational phenotyping methods.
[…]
@@
Hardware Requirements Minimum hardware requirements required to run model(s) of interest if considering internal environments. May be re-evaluated to accommodate any inference optimization strategies implemented[chitty-venkata_survey_2023]. Assess the maximum computational resources in each of the available host environments, focusing on graphical processing unit (GPU)resources in particular. Assess whether maximum resources will be available throughout the project or if a smaller model size is desired to ensure consistent access.
@@
Accuracy How accurate the model is at classifying or phenotyping based on a ground truth. Classification accuracy can be measured using common classification metrics like accuracy, precision, recall, and area under the receiver operating characteristic (AUROC).
@@
Time for Manual Review An estimated $66$ months of the total $1818$ months for ARF phenotype development were required for manual review of concepts. An week was required for clinical experts to collate information on the concepts of interest to be used for prompt engineering and an additional day was required to manually review model ability results. LLM-based method was superior.

### v3

@@ Conclusions
We developed PHEONA, an evaluation framework for applying LLM-based methods to computational phenotyping tasks. This framework helps assess if LLMs are feasible and if so, which model is best for each task. We demonstrated PHEONA on an example use case to promote further research into improving computational phenotyping.
@@
Hardware Requirements Minimum hardware requirements required to run model(s) of interest if considering internal environments. May be re-evaluated to accommodate any inference optimization strategies implemented[chitty-venkata_survey_2023]. Assess the maximum computational resources in each of the available host environments, focusing on graphics processing unit (GPU)resources in particular. Assess whether maximum resources will be available throughout the project or if a smaller model size is desired to ensure consistent access.
@@
Accuracy How accurate the model is at classifying or phenotyping based on a ground truth. Based on task requirements, this may also include assessing the accuracy of responses with respect to the reference text or provided context. Classification accuracy can be measured using common classification metrics like accuracy, precision, recall, and area under the receiver operating characteristic (AUROC). Accuracy against reference text can be performed using N-gram analysis, BLEU, or ROUGE while accuracy against provided context can be performed also by N-gram analysis and additionally ROUGE-C.[sai_survey_2022]
@@
Time for Manual Review An estimated $66$ months of the total $1818$ months for ARF phenotype development were required for manual review of concepts. A week was required for clinical experts to collate information on the concepts of interest to be used for prompt engineering and an additional day was required to manually review model ability results. LLM-based method was superior.